I need help with GDP calculations

Consider the following data on U.S. GDP:

Year Nominal GDP GDP Deflator
(in billions of dollars) (base year 1996)
_____________________________________________________________________________

2000 9,873 118
1999 9, 269 113


a. What was the growth rate of nominal GDP between 1999 and 2000? (Note: The growth rate is the percentage change from one period to the next.)

b. What was the growth rate of the GDP deflator between 1999 and 2000?

c. What was the real GDP in 1999 measured in 1996 prices?

d. What was real GDP in 2000measured in 1996 prices?

e. What was the growth rate of real GDP between 1999 and 2000?

f. Was the growth rate of nominal GDP higher or lower than the growth rate of real GDP?

Explain
Answer #1
Consider the following data on U.S. GDP:

Year Nominal GDP GDP Deflator
(in billions of dollars) (base year 1996)
_____________________________________________________________________________

2000 9,873 118
1999 9, 269 113
Yr Nominal GDP Real GDP GDP deflator
1996=100
1999 9269113 8202754.867 113
2000 9873118 8367049.153 118


a. What was the growth rate of nominal GDP between 1999 and 2000? (Note: The growth rate is the percentage change from one period to the next.)

Growth rate of Nominal GDP from 1999 to 2000
Nominal GDP growth rate= [(NGDP 2000 - NGDP 1999)/ NGDP 1999] x 100
                                          = [(9873118 - 9269113)/ 9269113] x 100
                                          = 0.065 x100
                                          = 6.5 %


b. What was the growth rate of the GDP deflator between 1999 and 2000?
growth rate of the GDP deflator = [(GDPdeflator 2000 - GDP deflator 1999)/ GDP deflator 1999] x 100
                                                 = [(118- 113)/113] x 100
                                         = 0.4424 x 100
                                         = 4.424 %


c. What was the real GDP in 1999 measured in 1996 prices?
Real GDP at base year 1996
Real GDP= nominal GDP/ GDP deflator in hundreths
GDP deflator in hundreths= 113/100 = 1.13
Real GDP in 1999 at 1996 prices = 9269113/1.13
                                                      = $ 8202754.867 billions


d. What was real GDP in 2000measured in 1996 prices?
Real GDP= nominal GDP/ GDP deflator in hundreths
GDP deflator in hundreths= 118/100 = 1.18
Real GDP in 2000 at 1996 prices = 9873118/1.18
                                                      = $ 8367049.153 billions


e. What was the growth rate of real GDP between 1999 and 2000?
Growth rate of real GDP from 1999 to 2000 is
Growth rate= [(RGDP 2000 - RGDP 1999)/ RGDP 1999] x 100
                                          = [(8367049.153 - 8202754.867 )/ 8202754.867 ] x 100
                                          = 0.0200 x100
                                          = 2 %


f. Was the growth rate of nominal GDP higher or lower than the growth rate of real GDP?

Explain
nominal GDP growth rate is higher than the real GDP growth rate. This is because real GDP is caliculated at base year price of 100, where as the nominal GDP is caliculated at the inflated prices which is above 100. We can observe that the GDP deflator rise from 113 to 118 from 1999 to 2000. this is due to the rise in the price level or inflation, which has given a higher growth rate to the nominal GDP when caliculated at market price.
For example: If a country produces a pizza for $10 in the base year 1996, if the GDPdiflato is 1.13 in 1999 the price of pizza is $11.3. and in 2000 the price of pizza is $11.8.
So, in nominal GDP we caliculate output at market prices, and in real GDP we caliculte output at base year price i.e $10.